# SLRQA: A Sparse Low-Rank Quaternion Model for Color Image Processing with Convergence Analysis

### v1

@@ 3 An Algorithm for SLRQA model
1:A quaternion matrix $\bm{\hat{{\mathbf{\dot{X}}}}}\in\mathbb{H}^{m\times n}$; a nonconvex function $\phi$; a constant $\mu>0$; a tolerance $\eta_{C}$; and an initial iterate $\bm{\sigma}_{0}=\mathbf{0}$;
2:$\bm{{\mathbf{\dot{X}}}}_{*}$;
@@
In Step 6, the $\mathbf{\dot{W}}$-subproblem approximately minimizes $\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})+\frac{L_{2,k}}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2}$ for a fixed ${\mathbf{\dot{X}}}_{k+1}$ and $\mathbf{\dot{\Lambda}}_{k}$. Specifically, the objective function of the $\mathbf{\dot{W}}$-subproblem is obtained by using a linearized strategy for $p(\mathbf{\dot{W}})$ in $\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})+\frac{L_{2,k}}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2}$. It follows that $\mathbf{\dot{W}}_{k+1}$ has a closed form solution
@@
$\mathbf{\dot{W}}_{k+1}=\frac{1}{\beta+L_{2,k}}\left(\beta\mathcal{W}({\mathbf{\dot{X}}}_{k+1})+\mathbf{\dot{\Lambda}}_{k}+L_{2,k}\mathbf{\dot{W}}_{k}\right)-\frac{\lambda}{\beta+L_{2,k}}\nabla p(\mathbf{\dot{W}}_{k}).$ (14)
@@
Step 7 is used to update the Lagrange multipliers. The difference between the classic ADMM for convex function and PL-ADMM is that $\mu\in(0,2)$ in PL-ADMM rather than $\mu\in(0,\frac{\sqrt{5}+1}{2})$ [8].
@@ 4 Convergence analysis
where $f$ is continuous and $g$ is continuously differentiable, $\mathcal{C},\mathcal{B}:\mathbb{H}^{m\times n}\rightarrow\mathbb{H}^{m\times n}$ are given linear maps and $\mathbf{\dot{B}}\in\mathbb{H}^{m\times n}$. Specifically, if $f({\mathbf{\dot{X}}})=\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma),g(\mathbf{\dot{W}})=p(\mathbf{\dot{W}})$, $\mathcal{B}=-\mathbf{I}$, $\mathcal{C}=\mathcal{W}$, and $\mathbf{\dot{B}}=0$, then Problem (15) becomes the SLRQA in (2).
The augmented Lagrangian function of (15) is
@@
The proposed PL-ADMM for (15) is
@@
$+\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta}\right\|_{\mathrm{F}}^{2}+\frac{L_{1,k}}{2}\|{\mathbf{\dot{X}}}-{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}^{2},$
$\mathbf{\dot{W}}_{k+1}$ $\in\mathop{\mathrm{arg\,min}}_{\mathbf{\dot{W}}}\left\langle\nabla g(\mathbf{\dot{W}}_{k}),\mathbf{\dot{W}}\right\rangle+\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta}\right\|_{\mathrm{F}}^{2}+\frac{L_{2,k}}{2}\left\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\right\|^{2}_{\mathrm{F}},$
@@
The choices of $\beta,L_{1,k}$ and $L_{2,k}$ are crucial for the convergence of (17). Generally speaking, $\beta$ should be chosen large enough, and $L_{1,k}$, $L_{2,k}$ should be chosen in a suitable range. In order to give the conditions of $\beta$ and $L_{1,k},L_{2,k}$, we introduce the following notations. Let $q_{i}:=\sup_{k\geq 0}{L_{i,k}}<\infty$ and $q_{i}^{-}:=\inf_{k>0}{L_{i,k}}>0$ for $i=1,2$. Define $\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}$ as the smallest positive eigenvalue of $\mathcal{B}^{\#}\mathcal{B}$, and $\rho(\mu):=1-|1-\mu|$. For any sequence $\{\mathbf{\dot{U}}_{k}\}_{k\geq 0}$, define $\Delta$ operator by $\Delta\mathbf{\dot{U}}_{k}:=\mathbf{\dot{U}}_{k}-\mathbf{\dot{U}}_{k-1}.$
The assumptions for the convergence analysis are given in Assumption 4.1, including the conditions of $\beta$, $L_{1,k}$, and $L_{2,k}$.
@@
    It has been shown in [2, Lemma 5.7] that (18) implies that for every $\mathbf{\dot{W}}_{1}$ and $\mathbf{\dot{W}}_{2}$, it holds that
@@
    The matrix $\mathcal{B}^{\#}\mathcal{B}$ is full rank, ${\rm range}(\mathcal{C})\subseteq{\rm range}(\mathcal{B})$, and $\mathbf{\dot{B}}\in{\rm range}(\mathcal{B})$.
@@
    $q_{1}^{-}\mathbf{I}\succeq a_{1}\mathbf{I}\,\quad\mbox{and}\,\quad q_{2}^{-}\mathbf{I}+\beta\mathcal{B}^{\#}\mathcal{B}-(r\theta_{0}+r\theta_{1}+L_{g})\mathbf{I}\succeq a_{2}\mathbf{I},$ (20)
@@
    $\theta_{0}:=\frac{2\mu(q_{2}+L_{g})^{2}}{\beta\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}(\rho(\mu))^{2}},\,\hbox{ and }\theta_{1}:=\frac{2\mu q_{2}^{2}}{\beta\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}(\rho(\mu))^{2}}.$ (21)
@@
    The parameter $\beta$ satisfies $\beta>\frac{2L_{g}}{\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}\rho(\mu)}.$
One can verify that A1, A2, and A3 in Assumption 4.1 are satisfied for SLRQA in (2). If $\beta$ is chosen sufficiently large and the sequences $\{L_{1,k}\}$ and $\{L_{2,k}\}$ are chosen such that $q_{i}^{-},q_{i}$, $i=1,2$ in suitable range, then A4 and A5 in Assumption 4.1 are also satisfied. For example, if set $L_{1,k}=L_{2,k}=1$ for all $k$ and let
@@
$\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})-\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\geq\frac{L_{1,k}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}.$
@@
$f({\mathbf{\dot{X}}}_{k+1})-f({\mathbf{\dot{X}}}_{k})+\frac{\beta}{2}\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta\|_{\mathrm{F}}^{2}+\frac{L_{1,k}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}$ (22)
@@
$-\frac{\beta}{2}\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta\|_{\mathrm{F}}^{2}+\frac{\beta}{2}\|\mathcal{C}({\mathbf{\dot{X}}}_{k})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta\|_{\mathrm{F}}^{2}\geq\frac{L_{1,k}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2},$
@@
where $B_{k}:=L_{2,k}\mathbf{I}-L_{g}\mathbf{I}+\frac{\beta}{2}\mathcal{B}\mathcal{B}^{\#}$.
@@
$\nabla g(\mathbf{\dot{W}}_{k})+\beta\mathcal{B}^{\#}(\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta)+L_{2,k}\Delta\mathbf{\dot{W}}_{k+1}=0.$ (23)
@@
$\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1})\leq\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})-\|\Delta\mathbf{\dot{W}}_{k+1}\|^{2}_{B_{k}}-\frac{L_{1,k}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}+\frac{1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}.$
@@
$\leq\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})-\|\Delta\mathbf{\dot{W}}_{k+1}\|_{B_{k}}^{2}-\frac{L_{1,k}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+\frac{1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}.$

### v2

@@ 3.1 An Algorithm for SLRQA
1:A quaternion matrix $\bm{\hat{{\mathbf{\dot{X}}}}}\in\mathbb{H}^{m\times n}$, a nonconvex function $\phi$, a constant $\mu>0$, a tolerance $\eta_{C}$, and an initial iterate $\bm{\sigma}_{0}=\mathbf{0}$.
2:$\bm{{\mathbf{\dot{X}}}}_{*}$.
@@
In Step 6, the $\mathbf{\dot{W}}$-subproblem approximately minimizes $\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})+\frac{L_{k,2}}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2}$ for a fixed ${\mathbf{\dot{X}}}_{k+1}$ and $\mathbf{\dot{\Lambda}}_{k}$. Specifically, the objective function of the $\mathbf{\dot{W}}$-subproblem is obtained by using a linearized strategy for $p(\mathbf{\dot{W}})$ in $\mathcal{L}_{\beta}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})+\frac{L_{k,2}}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2}$. It follows that $\mathbf{\dot{W}}_{k+1}$ has a closed form solution
@@
$\mathbf{\dot{W}}_{k+1}=\frac{1}{\beta+L_{k,2}}\left(\beta\mathcal{W}({\mathbf{\dot{X}}}_{k+1})+\mathbf{\dot{\Lambda}}_{k}+L_{k,2}\mathbf{\dot{W}}_{k}\right)-\frac{\lambda}{\beta+L_{k,2}}\nabla p(\mathbf{\dot{W}}_{k}).$ (19)
@@
Step 7 is used to update the Lagrange multiplier. The difference between the classic ADMM for convex function and PL-ADMM is that $\mu\in(0,2)$ in PL-ADMM rather than $\mu\in(0,\frac{\sqrt{5}+1}{2})$ [10], which means that a larger stepsize can be chosen with convergence guarantee.
3.2 An Algorithm for SLRQA-NF
The proposed algorithm PL-ADMM-NF for solving SLRQA-NF is stated in Algorithm 3.
Algorithm 3 PL-ADMM-NF algorithm for SLRQA-NF in (4)
1:Corrupted images $\mathbf{\dot{Y}}$, two parameters $\mu\in(0,2)$ and $\beta>0$, a tolerance $\eta>0$, a sequence of regularization parameters $\{L_{k,1}\},\{L_{k,2}\}$ satisfying $\sup_{k\geq 0}{L_{k,1}}<\infty$, $\sup_{k\geq 0}{L_{k,2}}<\infty$ and $\inf_{k>0}{L_{k,1}}>0$, $\inf_{k>0}{L_{k,2}}>0$.
2:Recovered $\mathbf{\dot{Z}}_{*}$.
4:for $k=0,1,2,\ldots$ do
5:     $\mathbf{\dot{Z}}$ subproblem:
$\mathbf{\dot{Z}}_{k+1}\in\mathop{\mathrm{arg\,min}}_{\mathbf{\dot{Z}}}\sum_{i}\phi(\sigma_{i}(\sqrt{{\mathbf{\dot{Z}}}^{*}\mathbf{\dot{Z}}+\varepsilon^{2}\mathbf{I}}),\gamma)+\left\langle\mathbf{\dot{\Lambda}}_{k,1},\mathbf{\dot{Z}}-\mathbf{\dot{W}}_{k}\right\rangle+\frac{\beta_{1}}{2}\|\mathbf{\dot{Z}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2}+\frac{L_{k,1}}{2}\|\mathbf{\dot{Z}}-\mathbf{\dot{Z}}_{k}\|_{\mathrm{F}}^{2};$
6:     $\mathbf{\dot{W}}$ subproblem:
$\mathbf{\dot{W}}_{k+1}$ $=\mathop{\mathrm{arg\,min}}_{\mathbf{\dot{W}}}\lambda\left\langle\nabla p(\mathbf{\dot{W}}_{k}),\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\right\rangle+\frac{\beta_{1}}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{Z}}_{k+1}+\frac{1}{\beta_{1}}\mathbf{\dot{\Lambda}}_{k,1}\|_{\mathrm{F}}^{2}$
$+\frac{\beta_{2}}{2}\left\|(\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}})-\mathbf{\dot{Y}})+\frac{1}{\beta_{2}}\mathbf{\dot{\Lambda}}_{k,2}\right\|_{\mathrm{F}}^{2}+\frac{L_{k,2}}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}^{2};$
7:     $\mathbf{\dot{\Lambda}}_{k+1,1}=\mathbf{\dot{\Lambda}}_{k,1}+\mu\beta(\mathbf{\dot{Z}}_{k+1}-\mathbf{\dot{W}}_{k+1})$;
8:     $\mathbf{\dot{\Lambda}}_{k+1,2}=\mathbf{\dot{\Lambda}}_{k,2}+\mu\beta\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{Y}})$;
10:         Return $\mathbf{\dot{Z}}_{*}=\mathbf{\dot{Z}}_{k+1}$; break
11:     end if
12:end for
The augmented Lagrangian function of SLRQA-NF (4) is given by:
~$\tilde{\mathcal{L}}_{\beta_{1},\beta_{2}}(\mathbf{\dot{Z}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{\Lambda}}_{2})=\sum_{i}\phi(\sigma_{i}(\sqrt{{\mathbf{\dot{Z}}}^{*}\mathbf{\dot{Z}}+\varepsilon^{2}\mathbf{I}}),\gamma)+\lambda p(\mathbf{\dot{W}})+\left\langle\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{Z}}-\mathbf{\dot{W}}\right\rangle$ (20)
$+\frac{\beta_{1}}{2}\|\mathbf{\dot{Z}}-\mathbf{\dot{W}}\|^{2}_{\mathrm{F}}+\left\langle\mathbf{\dot{\Lambda}}_{2},\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\right\rangle+\frac{\beta_{2}}{2}\|\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}$
$=\sum_{i}\phi(\sigma_{i}(\sqrt{{\mathbf{\dot{Z}}}^{*}\mathbf{\dot{Z}}+\varepsilon^{2}\mathbf{I}}),\gamma)+\lambda p(\mathbf{\dot{W}})+\frac{\beta_{1}}{2}\|\mathbf{\dot{Z}}-\mathbf{\dot{W}}+\frac{1}{\beta_{1}}\mathbf{\dot{\Lambda}}_{1}\|_{\mathrm{F}}^{2}$
$+\frac{\beta_{2}}{2}\|\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})+\frac{1}{\beta_{2}}\mathbf{\dot{\Lambda}}_{2}\|_{\mathrm{F}}^{2}-\frac{1}{2\beta_{1}}\|\mathbf{\dot{\Lambda}}_{1}\|_{\mathrm{F}}^{2}-\frac{1}{2\beta_{2}}\|\mathbf{\dot{\Lambda}}_{2}\|^{2}_{\mathrm{F}},$
where $\beta_{1},\beta_{2}>0$ are the penalty parameters, $\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{\Lambda}}_{2}\in\mathbb{H}^{m\times n}$ are the Lagrange multipliers.
In Step 3 of Algorithm 3, since the subproblem of $\mathbf{\dot{Z}}$ is equivalent to
$\min_{\mathbf{\dot{Z}}}\,\sum_{i}\phi(\sigma_{i}(\sqrt{{\mathbf{\dot{Z}}}^{*}\mathbf{\dot{Z}}+\varepsilon^{2}\mathbf{I}}),\gamma)+\frac{\beta+L_{k,1}}{2}\left\|\mathbf{\dot{Z}}-\frac{1}{\beta+L_{k,1}}\left(\beta\mathbf{\dot{W}}_{k}+\mathbf{\dot{\Lambda}}_{k,1}+L_{k,1}\mathbf{\dot{Z}}_{k}\right)\right\|_{\mathrm{F}}^{2},$
Algorithm 2 can be used to obtain $\mathbf{\dot{Z}}_{k+1}$. In Step 6 of Algorithm 3, the solution of $\mathbf{\dot{W}}$ has an explicit solution:
$\mathbf{\dot{W}}_{k+1}=\mathcal{G}^{-1}\left(\beta_{1}\mathbf{\dot{Z}}_{k+1}-\mathbf{\dot{\Lambda}}_{k,1}+\mathcal{W}\mathcal{A}^{\#}\mathcal{P}_{\Omega}^{\#}\left(\beta_{2}\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})-\mathbf{\dot{\Lambda}}_{k,2}\right)+L_{k,2}\mathbf{\dot{W}}_{k}-\lambda\nabla p(\mathbf{\dot{W}}_{k})\right),$
where $\mathcal{G}=(\beta_{1}+L_{k,2})\mathcal{I}+\beta_{2}\mathcal{W}\mathcal{A}^{\#}\mathcal{P}_{\Omega}^{\#}\mathcal{P}_{\Omega}\mathcal{A}\mathcal{W}^{\#}$. Steps 7 and 8 are used to update the Lagrange multipliers. Finally, the recovered image ${\mathbf{\dot{X}}}_{*}$ is obtained by ${\mathbf{\dot{X}}}_{*}=\mathcal{W}^{\#}(\mathbf{\dot{Z}}_{*})$.
@@ 4 Convergence analysis
In this section, we present the global convergence analysis for Algorithm 1 and 3 respectively. The convergence analysis for 1 is similar to that in [43]. However, the algorithm we analysis is slightly different from [43]. Furthermore, we present more flexible assumptions. The convergence analysis for Algorithm 3 is different from the previous works since two linear constraints are considered. To the best of our knowledge, this is the first convergence analysis for ADMM with two linear constrants.
4.1 Global Convergence for Algorithm 1
@@
where $f$ is semi-lowercontinuous and $g$ is continuously differentiable, $\mathcal{C},\mathcal{B}:\mathbb{H}^{m\times n}\rightarrow\mathbb{H}^{m\times n}$ are given linear maps and $\mathbf{\dot{B}}\in\mathbb{H}^{m\times n}$. Specifically, if $f({\mathbf{\dot{X}}})=\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma),g(\mathbf{\dot{W}})=p(\mathbf{\dot{W}})$, $\mathcal{B}=-\mathbf{I}$, $\mathcal{C}=\mathcal{W}$, and $\mathbf{\dot{B}}=0$, then Problem (21) becomes the SLRQA in (2). The augmented Lagrangian function of (21) is
@@
The iteration scheme for the proposed PL-ADMM for (21) is
@@
$+\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta}\right\|_{\mathrm{F}}^{2}+\frac{L_{k,1}}{2}\|{\mathbf{\dot{X}}}-{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}^{2},$
$\mathbf{\dot{W}}_{k+1}$ $\in\mathop{\mathrm{arg\,min}}_{\mathbf{\dot{W}}}\left\langle\nabla g(\mathbf{\dot{W}}_{k}),\mathbf{\dot{W}}\right\rangle+\frac{\beta}{2}\left\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}})-\mathbf{\dot{B}}+\frac{\mathbf{\dot{\Lambda}}_{k}}{\beta}\right\|_{\mathrm{F}}^{2}+\frac{L_{k,2}}{2}\left\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\right\|^{2}_{\mathrm{F}},$
@@
The choices of $\beta,L_{k,1}$ and $L_{k,2}$ are crucial for the convergence of (59). Generally speaking, $\beta$ should be chosen large enough, and $L_{k,1}$, $L_{k,2}$ should be chosen in a suitable range. In order to give the conditions of $\beta$ and $L_{k,1},L_{k,2}$, we introduce the following notations. Let $q_{i}:=\sup_{k\geq 0}{L_{i,k}}<\infty$ and $q_{i}^{-}:=\inf_{k>0}{L_{i,k}}>0$ for $i=1,2$. Define $\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}$ as the smallest positive eigenvalue of $\mathcal{B}^{\#}\mathcal{B}$, and $\rho(\mu):=1-|1-\mu|$. For any sequence $\{\mathbf{\dot{U}}_{k}\}_{k\geq 0}$, define $\Delta$ operator by $\Delta\mathbf{\dot{U}}_{k}:=\mathbf{\dot{U}}_{k}-\mathbf{\dot{U}}_{k-1}.$
The assumptions for the convergence analysis are given in Assumption 4.1, including the conditions of $\beta$, $L_{k,1}$, and $L_{k,2}$.
@@
    It has been shown in [4, Lemma 5.7] that (24) implies that for every $\mathbf{\dot{W}}_{1}$ and $\mathbf{\dot{W}}_{2}$, it holds that
@@
    The matrix $\mathcal{B}^{\#}\mathcal{B}$ is full rank or $g(\mathbf{\dot{W}})$ is coercive, ${\rm range}(\mathcal{C})\subseteq{\rm range}(\mathcal{B})$, and $\mathbf{\dot{B}}\in{\rm range}(\mathcal{B})$.
@@
    $q_{1}^{-}\mathbf{I}\succeq a_{1}\mathbf{I}\,\quad\mbox{and}\,\quad q_{2}^{-}\mathbf{I}+\beta\mathcal{B}^{\#}\mathcal{B}-(r\varsigma_{0}+r\varsigma_{1}+L_{g})\mathbf{I}\succeq a_{2}\mathbf{I},$ (26)
@@
    $\varsigma_{0}:=\frac{2\mu(q_{2}+L_{g})^{2}}{\beta\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}(\rho(\mu))^{2}},\,\hbox{ and }\varsigma_{1}:=\frac{2\mu q_{2}^{2}}{\beta\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}(\rho(\mu))^{2}}.$ (27)
@@
    The parameter $\beta$ satisfies $\beta>\frac{2L_{g}}{\kappa\lambda_{+}^{\mathcal{B}^{\#}\mathcal{B}}\rho(\mu)}$, where $\kappa\in(0,1)$ is a given constant.
One can verify that A1, A2, and A3 in Assumption 4.1 are satisfied for SLRQA in (2). If $\beta$ is chosen sufficiently large and the sequences $\{L_{k,1}\}$ and $\{L_{k,2}\}$ are chosen such that $q_{i}^{-},q_{i}$, $i=1,2$ in suitable range, then A4 and A5 in Assumption 4.1 are also satisfied. For example, if set $L_{k,1}=L_{k,2}=1$ for all $k$ and let
@@
$\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})-\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})\geq\frac{L_{k,1}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}.$
@@
$f({\mathbf{\dot{X}}}_{k+1})-f({\mathbf{\dot{X}}}_{k})+\frac{\beta}{2}\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta\|_{\mathrm{F}}^{2}+\frac{L_{k,1}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}$ (28)
@@
$-\frac{\beta}{2}\|\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta\|_{\mathrm{F}}^{2}+\frac{\beta}{2}\|\mathcal{C}({\mathbf{\dot{X}}}_{k})+\mathcal{B}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta\|_{\mathrm{F}}^{2}\geq\frac{L_{k,1}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2},$
@@
where $B_{k}:=L_{k,2}\mathbf{I}-L_{g}\mathbf{I}+\frac{\beta}{2}\mathcal{B}\mathcal{B}^{\#}$.
@@
$\nabla g(\mathbf{\dot{W}}_{k})+\beta\mathcal{B}^{\#}(\mathcal{C}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}+\mathbf{\dot{\Lambda}}_{k}/\beta)+L_{k,2}\Delta\mathbf{\dot{W}}_{k+1}=0.$ (29)
@@
$\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k+1},\mathbf{\dot{W}}_{k+1},\mathbf{\dot{\Lambda}}_{k+1})\leq\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})-\|\Delta\mathbf{\dot{W}}_{k+1}\|^{2}_{B_{k}}-\frac{L_{k,1}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|^{2}_{\mathrm{F}}+\frac{1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}.$
@@
$\leq\mathcal{L}_{\beta}({\mathbf{\dot{X}}}_{k},\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{k})-\|\Delta\mathbf{\dot{W}}_{k+1}\|_{B_{k}}^{2}-\frac{L_{k,1}}{2}\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}^{2}+\frac{1}{\beta\mu}\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|^{2}_{\mathrm{F}}.$